OpenStudy (anonymous):

find the value of b and c such that the function f(x) is continuous on the entire real line f(x) = x+1 , 1= 1

OpenStudy (anonymous):

I know how to do it if it was only one constant that I needed to find but with 2 where do I get the second equation from and how do I know which condition to apply ?

OpenStudy (perl):

you have two limits you can plug in $$ \large{ \lim_{x \to 1} ~x+1 = \lim_{x \to 1} ~x^2 +bx + c \\ \lim_{x \to 3} ~x+1 = \lim_{x \to 3} ~x^2 +bx + c }$$

OpenStudy (michele_laino):

we have to apply these condions: at x= 1: \[f\left( 1 \right) = 1 + 1 = 2 = 1 + b + c\] whereas at x=3: \[f\left( 3 \right) = 3 + 1 = 4 = 9 + 3b + c\]

OpenStudy (anonymous):

but why is it okay to use x=1 for the x2+bx+c part ?